# From Insights to Actions: The Impact of Interpretability and Analysis Research on NLP

### v1

@@ Abstract
Interpretability and analysis (IA) research is a growing subfield within NLP with the goal of developing a deeper understanding of the behavior or inner workings of NLP systems and methods. Despite growing interest in the subfield, a commonly voiced criticism is that it lacks actionable insights and therefore has little impact on NLP. In this paper, we seek to quantify the impact of IA research on the broader field of NLP. We approach this with a mixed-methods analysis of: (1) a citation graph of 185K+ papers built from all papers published at ACL and EMNLP conferences from 2018 to 2023, and (2) a survey of 138 members of the NLP community. Our quantitative results show that IA work is well-cited outside of IA, and central in the NLP citation graph. Through qualitative analysis of survey responses and manual annotation of 556 papers, we find that NLP researchers build on findings from IA work and perceive it is important for progress in NLP, multiple subfields, and rely on its findings and terminology for their own work. Many novel methods are proposed based on IA findings and highly influenced by them, but highly influential non-IA work cites IA findings without being driven by them. We end by summarizing what is missing in IA work today and provide a call to action, to pave the way for a more impactful future of IA research.
@@
Marius Mosbach1,2  Vagrant Gautamthanks: Authors contributed equally.3  Tomás Vergara-Brownefootnotemark: 4 Dietrich Klakow3  Mor Geva5 1Mila Quebec AI Institute    2McGill University    3Saarland University 4Pontificia Universidad Católica de Chile    5Tel Aviv University marius.mosbach@mila.quebecmorgeva@tauex.tau.ac.il
@@ 1 Introduction
The rapid progress made in the development of large language models (LLMs, Devlin et al. (2019); Radford et al. (2019); Raffel et al. (2020); Bommasani et al. (2022); Touvron et al. (2023); OpenAI et al. (2024); Team et al. (2024)) has had a profound impact on the field of natural language processing (NLP) (Gururaja et al., 2023). While these models demonstrate unprecedented performance and novel capabilities (Brown et al., 2020; Wei et al., 2022), and are rapidly finding their way into real-world applications (OpenAI, 2022; Microsoft, 2023; Google, 2024), they are also opaque and largely treated as black boxes, which does not satisfy other expectations for successful machine learning deployment, such as fairness, trust, accountability, and explainability (Lipton, 2018; Goodman and Flaxman, 2017).
@@
In NLP research, these factors have motivated a large body of work on interpretability and analysis (IA), which aims to understand the inner workings of LLMs and explain their predictions (Belinkov and Glass, 2019; Rogers et al., 2020; Rauker et al., 2023, inter alia). Researchers in this area are often motivated by the idea that better understanding LLMs is imperative to improve their efficiency, robustness, and trustworthiness, towards successful and safe deployment. IA research has thus witnessed rapid growth in the past few years and is now one of the biggest research areas (in terms of number of publications and citations) at the major NLP conferences (see Figure 1).
Despite the rapid growth of IA research (see also Figure 9), a commonly voiced criticism is that it often lacks actionable insights, especially for how to improve models, and therefore has little impact on how new NLP models are designed and built. This criticism raises questions about the usefulness of IA research, and whether its current form is the right path towards progress in NLP.
@@
While our findings show that IA work presents insightful observations, there are still opportunities for greater impact on the rest of NLP. Thus, based on survey responses, we identify the key ingredients that are missing in IA research today — unification; actionable recommendations; human-centered, interdisciplinary work; and standardized, robust methods — and close with a call to action with recommendations (§7). We hope our work paves the way towards a more impactful future for IA research as the field continues to grow.
@@ 2.1 Interpretability and analysis (IA) research
Interpretability research has a long tradition in Machine Learning as well adjacent fields like NLP (Tishby and Zaslavsky, 2015; Karpathy et al., 2015; Kim et al., 2018, inter alia). There is no single agreed upon definition of the term interpretability (see Lipton (2018) for a critical discussion), but two prominent types of interpretability research focus on post-hoc explainability or increasing the transparency of machine learning methods and models (Lipton, 2018; Madsen et al., 2024). Analysis research is an even broader term and one might argue that nearly every scientific paper contains some form of analysis. In NLP, however, many interpretability and analysis papers have in common that their primary contribution is an analysis that aims to advance our understanding of NLP in some way, e.g., by analyzing methods, models, or algorithms (Belinkov and Glass, 2019; Rogers et al., 2020).
@@ 2.2 Measuring impact
Our goal is to measure the impact of IA work on NLP research, which is not trivial to define, let alone quantify. To get a holistic view of impact, we consider two different, complementary ways of measuring impact – a bibliometric analysis, and a survey of the NLP community.
@@
Figure 2: Diagram showing the process of constructing our citation graph. Starting from an initial set of ACL and EMNLP papers we collect citations via the Semantic Scholar API and label papers with a classifier.
@@ 3.1 Citation graph construction
Figure 2 illustrates the process of constructing our citation graph. We start from an initial set of all papers published at ACL and EMNLP from 2018 to 2023. We focus on these two venues as they are leading NLP conferences with a dedicated track for interpretability and analysis research since 2020.222We discuss this decision in more detail in Limitations. Using this initial set of papers, we build a citation graph using the Semantic Scholar API Kinney et al. (2023). For papers outside our initial set, where we have gold labels, we rely on classifiers to predict submission tracks. More details on all these stages are provided below.
@@
We collect paper lists and track information from various sources (see Table 3 in Appendix B), as there is no one source of this data for ACL and EMNLP conferences. [^3] The ACL Anthology does not contain information on the submission track. Between 2018 and 2023, official names of submission tracks have changed substantially, so we standardize all data to 27 tracks. More details on this process are provided in Appendix B, including summary statistics per track (Table 1).
@@
We collect the citations of each paper in our initial set via the Semantic Scholar API (Kinney et al., 2023), resulting in a citation graph of 185,384 papers (see Table 2 in Appendix B for additional statistics). For each node (paper) in the graph, we store its title, abstract, and venue. For each edge (citation), we store information on the citation intent (binary labels for background, use of methods or comparing results), and citation influence (normal vs. highly influential), all of which are provided by Semantic Scholar.
@@
To assign all papers in the citation graph to our standardized set of tracks, we train a classifier based on the titles and abstracts from our initial set of papers. We find that some tracks are very hard to predict due to limited training data and the inherent ambiguity of submission tracks. We thus keep 11 well-performing labels (including IA), and introduce an ‘Other’ label to group the remaining papers. More details on classifier construction are provided in Appendix B.
Our final classifier achieves a test micro/macro-F1 score of 0.61/0.61. Although this performance appears rather low, we note that submission tracks have fuzzy boundaries, so papers can often be plausible submissions to multiple tracks. Given that we care primarily about accurately predicting IA compared to other tracks, we evaluate our classifier on two additional gold sets of data (see Section B.1) and obtain 78.1% and 87.8% accuracy on each set.
@@ 3.2 Surveying the NLP community
To solicit opinions from the NLP community on the impact of IA research, we ran a survey from March 19th to June 7th, 2024, advertising within our networks, on social media, and on NLP mailing lists. The full survey is shown in Appendix C.
Figure 3: Interpretability and analysis track CSI scores matrix against other tracks. These represent the probability that a random interpretability and analysis paper published in certain year has more citations than a random paper of other track published the same year.
To strike a balance between easy scoring and respondent expressivity, we included multiple-choice as well as optional free response questions (Shaughnessy et al., 2015). We refined the survey following best practices [^4] We made sure to clarify definitions, avoid leading questions, etc. (Shaughnessy et al., 2015). and with feedback from four senior NLP researchers who filled out a pilot version. We received a total of 138 responses from NLP researchers in academia and practitioners in industry, with 61% of respondents not working on IA themselves (see Appendix C for more statistics).
Two authors performed qualitative coding, an inductive method from the social sciences (Saldana, 2021), to identify themes in answers to the free-response questions. More details on the coding process are provided in Appendix D. We measure inter-coder reliability with percentage agreement (O’Connor and Joffe, 2020), which was above 90% across all subsets of annotation.
@@ 4 Researchers build on findings from IA research in their work
Figure 3 shows that CSI scores for the IA track are often favorable ($\text{CSI}>50\%$) when compared to other tracks. In 2023, only the Ethics and the Large Language Models tracks had favorable CSI scores against IA. This shows that IA papers have higher citational impact than other tracks, particularly in recent conferences.
@@
Figure 4: Origin of citations to IA papers.
@@
Figure 4 shows that most citations to IA papers are predicted to be extra-track citations. The proportion of references to IA papers differs considerably by citing track, with papers about Efficient Methods, Machine Learning, and Large Language Models citing IA research more frequently than others (see Figure 11 for a visualization of this trend). While the IA track does not stand out in terms of its extra-track citations compared to other tracks (see Figure 12), these results still demonstrate that the citational impact of IA research extends well beyond the IA track itself.
@@
Next, we assess whether IA papers are impacting NLP as a whole rather than just specific tracks. We quantify this with the Betweenness Centrality (BC) metric, a measure of interdisciplinarity Leydesdorff (2007); Barnett et al. (2011); Leydesdorff et al. (2018). BC quantifies the extent to which a node in the graph acts as a bridge along the shortest path between two other nodes Golbeck (2015); nodes with higher BC are considered more important as more information passes through them. [^5] We provide further discussion of BC in Section B.1. Therefore, we interpret papers with a high BC as important papers that are essential for the connectivity of the citation network.
@@
Figure 5: Survey responses on the frequency of using concepts from IA research, split by whether the respondents work in this field or not. Higher values indicate more frequent usage.
@@ 5 Researchers find IA work important
Figure 6: Survey responses (N=$138$) on whether progress in NLP in the last 5 years would have been slower or impossible without findings from interpretability and analysis research.
@@
When asked whether they thought IA work was important and if so, why, respondents overwhelmingly (133/138) consider it important, citing a variety of reasons, the most popular of which were: understanding model limitations and capabilities (90% of respondents), explainability for users (66%), improving model trustworthiness (59%), and improving model capabilities (50%). While a small percentage (4.3%) of respondents indicated that they thought it was not important (possibly also due to selection bias in our survey), we found that they voice the same concerns as those who do find it important, e.g., a lack of actionability, results that don’t scale, and a lack of impact on the most capable models of today. In our recommendations for the future of the field (§7), we go into these in more detail.
Figure 7: Survey responses (N=138) on how important interpretability and analysis research is to work in different subfields.
@@ 6 A closer look at influential papers
To this end, we inductively obtain the themes of a total of 585 papers, through qualitative coding of their titles and abstracts by two authors (Saldana, 2021). The 585 papers include: (1) All papers mentioned more than once as having influenced survey respondents’ work (N=29); (2) highly-cited IA papers from our citation graph (N=50); (3) highly-cited non-IA papers from our citation graph (N=50); (4) non-IA papers that cite and are highly influenced by the top-10 most-cited IA papers (N=456). The resulting themes are mostly descriptive, including topics (e.g., in-context learning, training dynamics) and contribution types (e.g., novel method, analysis). Percentage agreement on our coded themes is above 90% for each subset of papers. See Appendix D for more details.
@@
In contrast, the top-50 most cited IA papers are more often about the analysis component of IA (40%). Novel methods (for analysis, evaluation, linguistics, probing) are proposed in 26% of papers, and evaluation is a main contribution of 32%. As expected, the most cited non-IA papers in our citation graph mostly consist of highly influential datasets, models, and methods, e.g., HotpotQA, BART, prefix-tuning (Yang et al., 2018; Lewis et al., 2020; Li and Liang, 2021). More top themes are shown with the percentage of papers in Table 5 in Appendix E.
@@
As citations can have a variety of reasons (Zhu et al., 2015; Tahamtan and Bornmann, 2019), we examine three types of citational intent – background, methods and results citations (see Figure 13 in Appendix E). Overall, we find that influential IA papers are cited most often as background citations, then as methods citations, and least frequently when comparing results. In comparison, highly cited papers that are not about IA tend to be cited most frequently for methods. This is expected, as many of these papers are about popular datasets and models, as described above.
@@
Despite the large number of background citations, however, there is plenty of work—including non-IA work—that is highly influenced (according to Semantic Scholar) by IA research. For a closer look at what these citing papers do, we analyze all 456 papers with a highly influential citations to one of the top 10 most-cited IA papers, and annotate their themes based on titles and abstracts.

### v2

@@ Abstract
Interpretability and analysis (IA) research is a growing subfield within NLP with the goal of developing a deeper understanding of the behavior or inner workings of NLP systems and methods. Despite growing interest in the subfield, a criticism of this work is that it lacks actionable insights and therefore has little impact on NLP. In this paper, we seek to quantify the impact of IA research on the broader field of NLP. We approach this with a mixed-methods analysis [^1] Code and data publicly available at https://github.com/mmarius/interpretability-impact of: (1) a citation graph of 185K+ papers built from all papers published at ACL and EMNLP conferences from 2018 to 2023, and their references and citations, and (2) a survey of 138 members of the NLP community. Our quantitative results show that IA work is well-cited outside of IA, and central in the NLP citation graph. Through qualitative analysis of survey responses and manual annotation of 556 papers, we find that NLP researchers build on findings from IA work and perceive it as important for progress in NLP, multiple subfields, and rely on its findings and terminology for their own work. Many novel methods are proposed based on IA findings and highly influenced by them, but highly influential non-IA work cites IA findings without being driven by them. We end by summarizing what is missing in IA work today and provide a call to action, to pave the way for a more impactful future of IA research.
@@
Marius Mosbach1,2  Vagrant Gautamthanks: Authors contributed equally.3  Tomás Vergara-Brownefootnotemark: 4,5 Dietrich Klakow3  Mor Geva6 1Mila Quebec AI Institute    2McGill University    3Saarland University 4Pontificia Universidad Católica de Chile    5CENIA    6Tel Aviv University marius.mosbach@mila.quebecmorgeva@tauex.tau.ac.il
@@ 1 Introduction
The rapid progress made in the development of large language models (LLMs, Devlin et al. (2019); Radford et al. (2019); Raffel et al. (2020); Bommasani et al. (2022); Touvron et al. (2023); OpenAI et al. (2024); Team et al. (2024)) has had a profound impact on the field of natural language processing (NLP) (Gururaja et al., 2023). While these models demonstrate unprecedented performance and novel capabilities (Brown et al., 2020; Wei et al., 2022), and are rapidly finding their way into real-world applications (OpenAI, 2022; Microsoft, 2023; Google, 2024), they are also largely treated as black boxes, which does not satisfy other expectations for successful machine learning deployment, such as trust, accountability, and explainability (Lipton, 2018; Goodman and Flaxman, 2017).
@@
In NLP research, these factors have motivated a large body of work on interpretability and analysis (IA), which aims to understand the inner workings of LLMs and explain their predictions (Belinkov and Glass, 2019; Rogers et al., 2020; Rauker et al., 2023, inter alia). Many researchers in this area believe that better understanding LLMs is imperative to improve their efficiency, robustness, and trustworthiness, towards successful, safe deployment. IA research has thus witnessed rapid growth in recent years and is now one of the biggest research areas (in terms of number of publications and citations) at major NLP conferences (see Figure 1).
Despite the rapid growth of IA research (see also Figure 9), a criticism of this work is that it often lacks actionable insights, especially for how to improve models, and therefore has little impact on how new NLP models are designed and built (Rauker et al., 2023; Rai et al., 2024). This criticism raises questions about whether its current form is the right path towards progress in NLP.
@@
While our findings show that IA work presents insightful observations, there are still opportunities for greater impact on the rest of NLP. Thus, based on survey responses, we identify the key ingredients that are missing in IA research today — unification; actionable recommendations; human-centered, interdisciplinary work; and standardized, robust methods — and close with a call to action and recommendations (§7). We hope our work paves the way towards a more impactful future for IA research as the field continues to grow.
@@ 2.1 Interpretability and analysis (IA) research
Interpretability research has a long tradition in machine learning as well adjacent fields like NLP (Tishby and Zaslavsky, 2015; Karpathy et al., 2015; Kim et al., 2018, inter alia). There is no single agreed upon definition of the term interpretability (see Lipton (2018) and Li et al. (2022) for a critical discussion), but two prominent types of interpretability research focus on post-hoc explainability or increasing the transparency of machine learning methods and models (Lipton, 2018; Madsen et al., 2024). Analysis research is an even broader term and one might argue that nearly every scientific paper contains some form of analysis. In NLP, however, many interpretability and analysis papers have in common that their primary contribution is an analysis that aims to advance our understanding of NLP in some way, e.g., by analyzing methods, models, or algorithms (Belinkov and Glass, 2019; Rogers et al., 2020).
@@ 2.2 Measuring impact
Our goal is to measure the impact of IA work on NLP research, which is not trivial to define or quantify. For a holistic view of impact, we consider two complementary ways of measuring it – a bibliometric analysis, and a survey of the NLP community.
@@
Figure 2: Diagram showing the process of constructing our citation graph. Starting from an initial set of ACL and EMNLP papers between 2018 and 2023, we collect their citations and references via the Semantic Scholar API and label the submission track of the papers with a classifier.
@@ 3.1 Citation graph construction
As Figure 2 illustrates, we begin constructing our citation graph from an initial set of all papers published at ACL and EMNLP from 2018 to 2023. We focus on these two venues as they are leading NLP conferences with a dedicated track for interpretability and analysis research since 2020.333We discuss this decision in more detail in Limitations. We then use the Semantic Scholar API Kinney et al. (2023) to get all citations and references of these initial papers, and add them to our citation graph. For papers outside our initial set (where we have gold labels), we use a classifier to predict their submission track. More details on all these stages are provided below.
@@
We collect paper lists and track information from various sources (see Table 3 in Appendix A), as there is no one source of this data for ACL and EMNLP conferences. Between 2018 and 2023, official names of submission tracks have changed substantially, so we standardize all data to 27 tracks. More details on this process are provided in Appendix A, including summary statistics per track (Table 1).
@@
We collect the citations and references of each paper in our initial set via the Semantic Scholar API (Kinney et al., 2023), resulting in a citation graph of 185,384 papers (see Table 2 in Appendix A for additional statistics). For each node (paper) in the graph, we store its title, abstract, and venue. For each edge (citation), we store information on the citation intent (binary labels for background, use of methods or comparing results), and citation influence (normal vs. highly influential), all of which are provided by Semantic Scholar.
@@
To assign all papers in the citation graph to our standardized set of tracks, we train a classifier based on the titles and abstracts from our initial set of papers. We find that some tracks are very hard to predict due to limited training data and the inherent ambiguity of submission tracks. We thus keep 11 well-performing labels (including IA), and introduce an ‘Other’ label to group the remaining papers.
Our final classifier achieves a test micro/macro-F1 score of 0.61/0.61. Although this appears low, we note that submission tracks have fuzzy boundaries and papers can often be plausible submissions to multiple tracks. Given that we care primarily about predicting IA compared to other tracks, we evaluate our classifier on two additional sets of gold data, and obtain 78.1% and 87.8% accuracy on each set. We provide further details on classifier construction and evaluation, and we verify our findings with only gold labeled papers in Appendix A.
@@ 3.2 Surveying the NLP community
To solicit opinions from the NLP community on the impact of IA research, we ran a survey from March 19th to June 7th, 2024, advertising within our networks, on social media, and on NLP mailing lists. The full survey is shown in Appendix B.
Figure 3: CSI scores for the interpretability and analysis track are favorable (> 50%) when compared to other tracks. The CSI score represents the probability that a random interpretability and analysis paper published in certain year has more citations than a random paper of other track published the same year.
To strike a balance between easy scoring and respondent expressivity, we included multiple-choice as well as optional free response questions (Shaughnessy et al., 2015). We refined the survey following best practices [^4] We made sure to clarify definitions, avoid leading questions, etc. (Shaughnessy et al., 2015). and with feedback from four senior NLP researchers who filled out a pilot version. We received a total of 138 responses from NLP researchers in academia and practitioners in industry, with 61% of respondents not working on IA themselves (see Appendix B for more statistics).
Two authors performed qualitative coding, an inductive method from the social sciences (Saldana, 2021), to identify themes in answers to the free-response questions. More details on the coding process are provided in Appendix C. We measure inter-coder reliability with percentage agreement (O’Connor and Joffe, 2020), which was above 90% across all subsets of annotation.
@@ 4 Researchers build on findings from IA research in their work
Figure 3 shows that CSI scores for the IA track are often favorable ($\text{CSI}>50\%$) when compared to other tracks. In 2023, only the Ethics and the Large Language Models tracks had favorable CSI scores when compared to IA. This shows that IA papers have higher citational impact than other tracks.
@@
Figure 4: Origin of citations to IA papers in our citation graph. More citations come from non-IA work than IA work, showing citational impact beyond the subfield.
@@
Figure 4 shows that most citations to IA papers are predicted to be extra-track citations. The proportion of references to IA papers differs considerably by citing track, with papers about Efficient Methods, Machine Learning, and Large Language Models citing IA research more frequently than others (see Figure 11). While the IA track does not stand out in terms of its extra-track citations compared to other tracks (see Figure 12), these results still demonstrate that the citational impact of IA research extends well beyond the IA track itself.
@@
Next, we assess whether IA papers are impacting NLP as a whole rather than just specific tracks. We quantify this with the Betweenness Centrality (BC) metric, a measure of interdisciplinarity Leydesdorff (2007); Barnett et al. (2011); Leydesdorff et al. (2018). BC quantifies the extent to which a node in the graph acts as a bridge along the shortest path between two other nodes Golbeck (2015); nodes with higher BC are considered more important as more information passes through them. [^5] We provide further discussion of BC in Section A.1. Therefore, we interpret papers with a high BC as important papers that are essential for the connectivity of the citation network.
@@
Figure 5: Survey responses on the frequency of using concepts from IA research, split by whether the respondents work in this field or not.
@@ 5 Researchers find IA work important
Figure 6: Survey responses (N=$138$) on whether progress in NLP in the last 5 years would have been slower or impossible without findings from interpretability and analysis research. Respondents believe progress in NLP would be slower but not impossible without IA.
@@
When asked whether they thought IA work was important and if so, why, respondents overwhelmingly (133/138) consider it important, citing a variety of reasons, the most popular of which were: understanding model limitations and capabilities (90% of respondents), explainability for users (66%), improving model trustworthiness (59%), and improving model capabilities (50%). While a small percentage (4.3%) of respondents indicated that they thought it was not important (possibly also due to selection bias in our survey), we found that they voice the same concerns as those who do find it important, e.g., a lack of actionability, results that do not scale, and a lack of impact on the most capable models of today. In our recommendations for the future of the field (§7), we go into these in more detail.
Figure 7: Survey responses (N=138) on how important interpretability and analysis research is to work in different subfields. IA is considered most important for work on reasoning, factuality, and bias, and least imporatnt for LLM engineering.
@@ 6 A closer look at influential papers
To this end, we inductively obtain the themes of a total of 585 papers, through qualitative coding of their titles and abstracts by two authors (Saldana, 2021). The 585 papers include: (1) All papers mentioned more than once as having influenced survey respondents’ work (N=29); (2) highly-cited IA papers from our citation graph (N=50); (3) highly-cited non-IA papers from our citation graph (N=50); (4) non-IA papers that cite and are highly influenced by the top-10 most-cited IA papers (N=456). The resulting themes are mostly descriptive, including topics (e.g., in-context learning, training dynamics) and contribution types (e.g., novel method, analysis). Percentage agreement on our coded themes is above 90% for each subset of papers. See Appendix C for more details.
@@
In contrast, the top-50 most cited IA papers are more often about the analysis component of IA (40%). Novel methods (for analysis, evaluation, linguistics, probing) are proposed in 26% of papers, and evaluation is a main contribution of 32%. As expected, the most cited non-IA papers in our citation graph mostly consist of highly influential datasets, models, and methods, e.g., HotpotQA, BART, prefix-tuning (Yang et al., 2018; Lewis et al., 2020; Li and Liang, 2021). More top themes are shown with the percentage of papers in Table 7 in Appendix D.
@@
As citations can have a variety of reasons (Zhu et al., 2015; Tahamtan and Bornmann, 2019), we examine three types of citational intent – background, methods and results citations (see Figure 13 in Appendix D). Overall, we find that influential IA papers are cited most often as background citations, then as methods citations, and least frequently when comparing results. In comparison, highly cited papers that are not about IA tend to be cited most frequently for methods. This is expected, as many of these papers are about popular datasets and models, as described above.
@@
Despite the large number of background citations, however, there is plenty of work—including non-IA work—that is highly influenced (according to Semantic Scholar) by IA research. For a closer look at what these citing papers do, we analyze all 456 papers with highly influential citations to one of the top 10 most-cited IA papers, and annotate their themes based on titles and abstracts.